The sum of two consecutive integers is 5/6 but the lower if the two is definitely negative so:
1/ x + 1/x+1 = 5/6
X is negative

Answer by stanbon(75887)

The sum of the inverse of two consecutive integers is 5/6 but the lower of the two is definitely negative so:
1/ x + 1/(x+1) = 5/6
X is negative
----
[x+1+x]/(x^2+x) = 5/6
------
12x+6 = 5x^2+5x
------
5x^2 - 7x - 6 = 0
------
5x^2 -10x+3x - 6 = 0
5x(x-2)+3(x-2) = 0
(x-2)(5x+3) = 0
---
Negative solution::
x = -3/5
x + 1 = 2/5
-------------
Note:: These are not integers.
Cheers,
Stan H.
--------------
----------------